# Near-Optimal Decision Trees in a SPLIT Second

### v2

@@ 5.2 LicketySPLIT: Polynomial-time SPLIT
Algorithm 3 $\ell$, $D$, $\lambda$, $d$)
0:  $\ell$, $D$, $\lambda$, $d$loss function, samples, regularizer, full depth}
0:  $t_{lookahead}=$ SPLIT$(\ell,D,\lambda,1,d,0)$Call SPLIT with lookahead depth $1$ and no post-processing}
@@ 5.3 RESPLIT: Rashomon set Estimation with SPLIT
At the cutting edge of compute requirements for decision tree optimization is the computation of Rashomon sets of decision trees. Xin et al. (2022) compute a Rashomon set of all near-optimal trees, based on the GOSDT algorithm (Lin et al., 2020). This task generates an extraordinary number of trees and has high memory and runtime costs. To make this tractable, Xin et al. (2022) leverage depth constraints and feature selection from prior work to reduce the depth and set of features considered (McTavish et al., 2022). While necessary for scalability, this can prevent exploration of near-optimal models across all features or at greater decision tree depths. Both factors are relevant for work on variable importance based on Rashomon sets (Fisher et al., 2019; Dong & Rudin, 2020; Donnelly et al., 2023). We leverage SPLIT as a way to dramatically improve scalability of Rashomon set computation, reliably approximating the full Rashomon set and allowing feasible exploration while relaxing or removing depth and feature constraints.
Our algorithm, RESPLIT, is described in Appendix A.10; it first leverages SPLIT as a subroutine to obtain a set of prefix trees such that completing them greedily up to the depth budget would result in an $\epsilon$ approximation of the optimal solution to Equation 5.1. At each leaf of each prefix tree, it calls TreeFARMS (Xin et al., 2022) to find a large set of shallow subtrees that are at least as good as being greedy, yielding an approximate Rashomon set computed much faster than state of the art. We also show a novel indexing mechanism to query RESPLIT trees in Appendix A.11.
Figure 3: Regularized test loss vs training time (in seconds) for GOSDT (McTavish et al., 2022) vs our algorithms. The size of the points indicates the number of leaves in the resulting tree. Both SPLIT and LicketySPLIT are much faster for most values of sparsity penalty $\lambda$, with the only potential slowdown being in the sub-second regime due to overhead costs.
@@ 7 Experiments
Figure 4: A comparison between the performance of our algorithms and competitors (depth budget $5$, lookahead depth $2$). The red box in the upper plot illustrates the region containing sparse and accurate models. The lower plots show the test loss vs training time for models in the red box. SPLIT and LicketySPLIT consistently lie on the bottom left of the test loss-sparsity frontier, with runtimes orders of magnitude faster than many competitors. Our algorithms also offer the ideal compromise between runtime and loss. All metrics are averaged over $3$ test-train splits.
@@ 7.1 How do our algorithms compare to GOSDT?
[…]
@@ 7.2 Characterising the Frontier of Test Loss, Sparsity, and Runtime
Figure 4 characterises the frontier of training time, sparsity, and test loss for several algorithms. Here, we vary hyper-parameters associated with each algorithm to produce trees of varying sparsity levels (where sparsity is the number of leaves). We see that there exists a frontier between test loss and sparsity, and different methods lie on different parts of the frontier. To maximize interpretability and accuracy, we want a tree to lie in the bottom left corner of the frontier, within the highlighted red rectangle. Out of all algorithms tested, ours consistently lie on the frontier and in the red rectangle. Alongside state of the art performance, our algorithms are often over 100$\times$ faster than their contemporaries. For more datasets, see Figure 5 in the Appendix.
@@ 7.3 Rashomon Set Approximation
We now show that RESPLIT enables fast, accurate approximation of the Rashomon set of near-optimal trees, while scaling much more favorably than state-of-the-art method TreeFARMS (Xin et al., 2022). We demonstrate that variable importance conclusions using RID (Donnelly et al., 2023) remain almost identical under RESPLIT, relative to the full Rashomon set. That is, RESPLIT allows accurate summary statistics of the full Rashomon set to be computed at greater depths and over more binary features while enhancing scalability. Table 1 shows computation of RID with and without RESPLIT. RESPLIT enables 1020×10-20$\times$10 - 20 × faster variable importance computation. Furthermore, the correlation between variable importances is very close to $1$, suggesting that RESPLIT trees serve as good proxies for estimating importances derived from the complete Rashomon set. Table 2 also shows that most of the trees output by RESPLIT lie in the true Rashomon set or very close to it.
@@
Table 1: Table summarizing the advantages of RESPLIT. The first $2$ columns show the time taken to compute all bootstrapped Rashomon sets for the Rashomon Importance Distribution (RID) (Donnelly et al., 2023) with and without RESPLIT. ##\## of bootstrapped datasets $=10$, $\lambda=0.02$, $\epsilon=0.01$, depth budget $5$, lookahead depth $3$. The last column shows the Pearson correlation between variable importances computed by RID and RID + RESPLIT. There is nearly perfect correlation seen in every case.
Dataset Precision Precision (Slack .$\mathbf{.01}$)
@@
Table 2: Proportion of RESPLIT Trees in the true Rashomon set (precision) and within at most .01.01.01.01 loss of being in the set. Most of the trees output by RESPLIT end up being in the Rashomon set. Trees which are not in the Rashomon set are almost always very close to being in it. We employ the same parameters as Table 1.
@@ Acknowledgements
We acknowledge funding from the National Institutes of Health under 5R01-DA054994, the National Science Foundation under award NSF 2147061, and through the Department of Energy under grant DE-SC0023194. We thank Srikar Katta for helpful discussions and feedback throughout this project.
@@ A.1.1 More datasets with depth $5$ trees
In Section 7 of the paper, we showed results for three datasets. Here, we evaluate SPLIT, LicketySPLIT, and its contemporaries on $6$ additional datasets. All datasets were evaluated on three random $80$-$20$ train-test splits of the data, with the average and standard error reported. Results are in Figure 5. Note that Covertype has smaller error bars because the dataset size is much larger – it has $\sim 5\times 10^{6}$ examples, while COMPAS and HELOC have only $\sim 10^{4}$ examples.
@@ A.1.2 What about depth $4$ trees?
Figure 6: A performance comparison between our algorithm and those in literature. Depth $4$ – Lookahead depth $2$.
@@ A.1.3 What about depth $6$ trees?
Figure 7: A performance comparison between our algorithm and those in literature. Depth $6$ – Lookahead depth $2$.
@@ A.2 Many Near-Optimal Trees Exhibit Monotonically Decreasing Optimality Gaps Closer to Leaves
$\beta(T,D,\ell)=\frac{\sum\limits_{\textrm{$t$}\in T}\delta(D_{t},t)\mathbbm{1}[\textrm{$t$ is rooted at level $\ell$}]}{{\sum\limits_{\textrm{$t$}\in T}\mathbbm{1}[\textrm{$t$ is rooted at level $\ell$}]}}.$ (6)
@@ A.3.2 Are SPLIT Trees in the Rashomon Set?
Figure 12: An illustration of near-optimality of our algorithms for depth budget $5$. The light yellow region represents the $(\lambda,\epsilon)$ configurations for which only SPLIT produces trees in the Rashomon set, while the darker region represents $(\lambda,\epsilon)$ values for which both SPLIT and LicketySPLIT produce trees in the Rashomon set. The figure shows that our trees are almost always in the Rashomon set even for small values of ($\epsilon$, $\lambda$).
@@ A.3.3 SPLIT with Optimality Preserving Discretization
  • We then produce two versions of the dataset – a fully binarized version (which contains around $3000$-$5000$ features for each dataset), and a threshold-guessed version (McTavish et al., 2022) with num__\__estimators $=200$. The latter ensures that the number of features in the resulting datasets is between $40$-$60$.
@@
We see that there is almost no difference in loss between the fully binarized dataset and the threshold guessed dataset, suggesting that there is minimial sacrifice in performance when using SPLIT / LicketySPLIT with threshold guessing. Furthermore, using threshold guessing results in runtimes that are orders of magnitude faster. These observations have also been corroborated by McTavish et al. (2022), though in the context of vanilla GOSDT.
@@
Let $T_{\mathcal{L},g}$ be the tree representing the solution to this equation - this is a lookahead prefix tree with greedy splits after depth $d_{l}$. Let $T_{SPLIT}$ be the tree that replaces the greedy subtree after depth $d_{l}$ with optimal GOSDT splits - this refers to lines $3$-$9$ in Algorithm 2. Let $T_{LSPLIT}$ be the tree that replaces the greedy subtree after depth $d_{l}$ with recursive LicketySPLIT subtrees (this refers to lines $3$-$7$ in Algorithm 3). We then vary the value of the sparsity penalty $\lambda\in[10^{-3},10^{-1}]$ and compute the post-processing gaps on the training dataset $D$:
@@
[…]
@@ A.4 SPLIT and LicketySPLIT Scaling Experiments
We now evaluate the scalability of SPLIT and its variants as the number of features increases. For each dataset evaluated, we use the threshold guessing mechanism from (McTavish et al., 2022) to binarize the dataset. In particular:
@@
In this experiment, we choose $n_{est}$ in a logarithmically spaced interval between $20$ and $10^{4}$, to obtain binary datasets with $10$-$1000$ features. We set a conservative value of $\lambda=0.001$ for SPLIT / LicketySPLIT, as from Figure 12, this ensures that the optimality gap for our method is around $\sim 10^{-3}$.
@@ A.5 Rashomon Importance Distribution Under RESPLIT vs TreeFARMS: Threshold Guessing
In this section, we compare RESPLIT and TreeFARMS in terms of their ability to generate meaningful variable importances under the Rashomon Importance Distribution (Donnelly et al., 2023). This analysis is a more complete representation of that in Table 1. The variable importance metric considered is Model Reliance (MR) - the precise details of how this is computed are in (Donnelly et al., 2023).
Figure 16: (top) Model Reliance for the top $5$ features when the Rashomon Importance Distribution is computed in its original form (with TreeFARMS), and when RESPLIT is used as the Rashomon set generating algorithm. The reported Pearson correlation is computed between the top $20$ features. We see that it is very close to $1$, i.e. features that are important under RID will also remain important when RESPLIT is used. (bottom) The number of models across the bootstrapped Rashomon sets which split on a given feature. We note from the bar plots that RESPLIT is also able to generate a large number of trees - often times as many as TreeFARMS.
Parameters: $\lambda=0.02,\epsilon=0.01,\#$ bootstrapped datasets $=10$, depth budget $=5$, lookahead depth $=3$.
@@ A.6 Rashomon Importance Distribution Under RESPLIT: Quantile Binarization
In this section, we show similar results as in the previous section, but when datasets are binarized using feature quantiles. We chose $3$ quantiles per feature (corresponding to each $3$rd of the distribution), resulting in datasets with 3×3$\times$3 × the number of features. For most of these datasets, RID with TreeFARMS failed to run in reasonable time.
Figure 17: (top) Model Reliance under RESPLIT for the top $5$ features. (bottom) The number of models across the bootstrapped Rashomon sets which split on a given feature. We note that the features which are important for RESPLIT under threshold guessing are also similarly important under quantile binarization, suggesting that our approach can generalize to different binarization schemes.
Parameters: $\lambda=0.02,\epsilon=0.01,\#$ bootstrapped datasets $=10$, depth budget $=5$, lookahead depth $=3$.
@@ A.7.1 Datasets
  • The Home Equity Line of Credit (HELOC) (FICO, 2018) dataset used for the Explainable ML Challenge. This dataset aims to predict the risk of loan default given the credit history of an individual. It consists of 23 features related to financial history, including FICO (credit) score, loan amount, number of delinquent accounts, credit inquiries, and other credit performance indicators. The dataset contains approximately 10,000 instances.
@@
  • Two recidivism datasets (COMPAS and Netherlands). COMPAS aims to predict the likelihood of recidivism (reoffending) for individuals who have been arrested. The dataset consists of approximately $6000$ instances and includes $11$ features including demographic attributes, criminal history, risk of general recidivism, and chargesheet information. The Netherlands dataset is a similar recidivism dataset containing demographic and prior offense features for individuals, used to predict reoffending risk.
@@
  • The Covertype dataset, which aims to predict the forest cover (one of $7$ types) for areas of the Roosevelt National Forest in northern Colorado, based on cartographic data. It contains $54$ attributes derived from US Geological Survey data. These include continuous variables like elevation, aspect, slope, and others related to soil type and climate. The dataset has over 580,000 instances, each corresponding to a $30$m $\times$ $30$m patch of the forest.
@@
  • The Adult dataset, which aims to predict whether an individual’s income exceeds \$$\$50000$ per year based on demographic and occupational information. It contains around $50000$ train and test examples, with $14$ features.
@@
  • The Bike dataset (Fanaee-T & Gama, 2013), which contains a two-year historical log of bikeshare counts from $2011$-$2012$ in Washington D.C., USA. It contains features relating to the weather at every hour – with the aim being to predict the number of bike rentals in the city in that given time period.
@@
  • The Hypothyroid dataset, which contains medical records used to predict whether a patient has hypothyroidism based on thyroid function test results and other medical attributes. It includes categorical and continuous variables such as TSH (thyroid-stimulating hormone) levels, age, and presence of goiter, with thousands of instances.
@@
  • The Spambase dataset, which consists of email data used to classify messages as spam or not spam. The dataset contains 57 features extracted from email text, such as word frequencies, capital letter usage, and special character counts, with around 4600 instances.
@@
  • The Bank dataset, which is used to predict whether a customer will subscribe to a bank term deposit based on features like age, job type, marital status, education level, and past marketing campaign success. It consists of approximately 4500 instances with 16 attributes.
@@
  • The HIV dataset contains RNA samples from $2$ patients. The labels correspond to whether the observed HIV viral load is high or not.

### v3

@@ 5.2 LicketySPLIT: Polynomial-time SPLIT
Algorithm 3 LicketySPLIT($\ell$, $D$, $\lambda$, $d$)
[…]
@@ 5.3 RESPLIT: Rashomon set Estimation with SPLIT
At the cutting edge of compute requirements for decision tree optimization is the computation of Rashomon sets of decision trees. xin2022treefarms compute a Rashomon set of all near-optimal trees, based on the GOSDT algorithm (gosdt). This task generates an extraordinary number of trees and has high memory and runtime costs. To make this tractable, xin2022treefarms leverage depth constraints and feature selection from prior work to reduce the depth and set of features considered (gosdt_guesses). While necessary for scalability, this can prevent exploration of near-optimal models across all features or at greater decision tree depths. Both factors are relevant for work on variable importance based on Rashomon sets (fisher2018model; dong2020exploring; donnelly2023the). We leverage SPLIT as a way to dramatically improve scalability of Rashomon set computation, reliably approximating the full Rashomon set and allowing feasible exploration while relaxing or removing depth and feature constraints.
Our algorithm, RESPLIT, is described in Appendix A.10; it first leverages SPLIT as a subroutine to obtain a set of prefix trees such that completing them greedily up to the depth budget would result in an $\epsilon$ approximation of the optimal solution to Equation 5.1. At each leaf of each prefix tree, it calls TreeFARMS (xin2022treefarms) to find a large set of shallow subtrees that are at least as good as being greedy, yielding an approximate Rashomon set computed much faster than state of the art. We also show a novel indexing mechanism to query RESPLIT trees in Appendix A.11.
Figure 3: Regularized test loss vs training time (in seconds) for GOSDT (gosdt_guesses) vs our algorithms. The size of the points indicates the number of leaves in the resulting tree. Both SPLIT and LicketySPLIT are much faster for most values of sparsity penalty $\lambda$, with the only potential slowdown being in the sub-second regime due to overhead costs.
@@ 7 Experiments
Figure 4: A comparison between the performance of our algorithms and competitors (depth budget $5$, lookahead depth $2$). The red box in the upper plot illustrates the region containing sparse and accurate models. The lower plots show the test loss vs training time for models in the red box. SPLIT and LicketySPLIT consistently lie on the bottom left of the test loss-sparsity frontier, with runtimes orders of magnitude faster than many competitors. Our algorithms also offer the ideal compromise between runtime and loss. All metrics are averaged over 33 test-train splits.
@@ 7.1 How do our algorithms compare to GOSDT?
  • When all methods have lower regularized objective values (left side of each plot), our methods are orders of magnitude faster than GOSDT. For instance, on the Bike dataset, SPLIT has training times of $\sim$10 seconds, while GOSDT runs for $\sim$$10^{3}$ seconds. LicketySPLIT takes merely a second in most cases. This is the regime most relevant to our algorithms.
@@ 7.2 Characterising the Frontier of Test Loss, Sparsity, and Runtime
[…]
@@ 7.3 Rashomon Set Approximation
We now show that RESPLIT enables fast, accurate approximation of the Rashomon set of near-optimal trees, while scaling much more favorably than state-of-the-art method TreeFARMS (xin2022treefarms). We demonstrate that variable importance conclusions using RID (donnelly2023the) remain almost identical under RESPLIT, relative to the full Rashomon set. That is, RESPLIT allows accurate summary statistics of the full Rashomon set to be computed at greater depths and over more binary features while enhancing scalability. Table 1 shows computation of RID with and without RESPLIT. RESPLIT enables $10-20\times$ faster variable importance computation. Furthermore, the correlation between variable importances is very close to $1$, suggesting that RESPLIT trees serve as good proxies for estimating importances derived from the complete Rashomon set. Table 2 also shows that most of the trees output by RESPLIT lie in the true Rashomon set or very close to it.
@@
Table 1: Table summarizing the advantages of RESPLIT. The first 22 columns show the time taken to compute all bootstrapped Rashomon sets for the Rashomon Importance Distribution (RID) (donnelly2023the) with and without RESPLIT. #\# of bootstrapped datasets =10=10, $\lambda=0.02$, $\epsilon=0.01$, depth budget $5$, lookahead depth $3$. The last column shows the Pearson correlation between variable importances computed by RID and RID + RESPLIT. There is nearly perfect correlation seen in every case.
Dataset Precision Precision (Slack $\mathbf{.01}$)
@@
Table 2: Proportion of RESPLIT Trees in the true Rashomon set (precision) and within at most .01.01 loss of being in the set. Most of the trees output by RESPLIT end up being in the Rashomon set. Trees which are not in the Rashomon set are almost always very close to being in it. We employ the same parameters as Table 1.
@@ Acknowledgements
We acknowledge funding from the National Institutes of Health under 5R01-DA054994, the National Science Foundation under award NSF 2147061, and through the Department of Energy under grant DE-SC0023194. We thank Srikar Katta, Jon Donnelly, Zachery Boner, Yixiao Wang, and Zakk Heile for helpful discussions and feedback throughout this project.
@@ A.1.1 More datasets with depth $5$ trees
In Section 7 of the paper, we showed results for three datasets. Here, we evaluate SPLIT, LicketySPLIT, and its contemporaries on $6$ additional datasets. All datasets were evaluated on three random 8080-2020 train-test splits of the data, with the average and standard error reported. Results are in Figure 5. Note that Covertype has smaller error bars because the dataset size is much larger – it has $\sim 5\times 10^{6}$ examples, while COMPAS and HELOC have only $\sim 10^{4}$ examples.
@@ A.1.2 What about depth $4$ trees?
Figure 6: A performance comparison between our algorithm and those in literature. Depth 44 – Lookahead depth $2$.
@@ A.1.3 What about depth $6$ trees?
Figure 7: A performance comparison between our algorithm and those in literature. Depth 66 – Lookahead depth $2$.
@@ A.2 Many Near-Optimal Trees Exhibit Monotonically Decreasing Optimality Gaps Closer to Leaves
[…]
@@ A.3.2 Are SPLIT Trees in the Rashomon Set?
Figure 12: An illustration of near-optimality of our algorithms for depth budget 55. The light yellow region represents the $(\lambda,\epsilon)$ configurations for which only SPLIT produces trees in the Rashomon set, while the darker region represents $(\lambda,\epsilon)$ values for which both SPLIT and LicketySPLIT produce trees in the Rashomon set. The figure shows that our trees are almost always in the Rashomon set even for small values of ($\epsilon$, $\lambda$).
@@ A.3.3 SPLIT with Optimality Preserving Discretization
  • We then produce two versions of the dataset – a fully binarized version (which contains around 30003000-$5000$ features for each dataset), and a threshold-guessed version (gosdt_guesses) with num_\_estimators =200=200. The latter ensures that the number of features in the resulting datasets is between 4040-$60$.
@@
We see that there is almost no difference in loss between the fully binarized dataset and the threshold guessed dataset, suggesting that there is minimial sacrifice in performance when using SPLIT / LicketySPLIT with threshold guessing. Furthermore, using threshold guessing results in runtimes that are orders of magnitude faster. These observations have also been corroborated by gosdt_guesses, though in the context of vanilla GOSDT.
@@
Let $T_{\mathcal{L},g}$ be the tree representing the solution to this equation - this is a lookahead prefix tree with greedy splits after depth $d_{l}$. Let $T_{SPLIT}$ be the tree that replaces the greedy subtree after depth $d_{l}$ with optimal GOSDT splits - this refers to lines 33-99 in Algorithm 2. Let $T_{LSPLIT}$ be the tree that replaces the greedy subtree after depth $d_{l}$ with recursive LicketySPLIT subtrees (this refers to lines 33-77 in Algorithm 3). We then vary the value of the sparsity penalty $\lambda\in[10^{-3},10^{-1}]$ and compute the post-processing gaps on the training dataset $D$:
@@
Figure 14: Gap (in $\%$ points) in accuracy between SPLIT / LicketySPLIT and a lookahead prefix tree followed by a purely greedy approach. Depth budget =5=5.
@@ A.4 SPLIT and LicketySPLIT Scaling Experiments
We now evaluate the scalability of SPLIT and its variants as the number of features increases. For each dataset evaluated, we use the threshold guessing mechanism from (gosdt_guesses) to binarize the dataset. In particular:
@@
In this experiment, we choose $n_{est}$ in a logarithmically spaced interval between $20$ and $10^{4}$, to obtain binary datasets with 1010-$1000$ features. We set a conservative value of $\lambda=0.001$ for SPLIT / LicketySPLIT, as from Figure 12, this ensures that the optimality gap for our method is around $\sim 10^{-3}$.
@@ A.5 Rashomon Importance Distribution Under RESPLIT vs TreeFARMS: Threshold Guessing
In this section, we compare RESPLIT and TreeFARMS in terms of their ability to generate meaningful variable importances under the Rashomon Importance Distribution (donnelly2023the). This analysis is a more complete representation of that in Table 1. The variable importance metric considered is Model Reliance (MR) - the precise details of how this is computed are in (donnelly2023the).
Figure 16: (top) Model Reliance for the top 55 features when the Rashomon Importance Distribution is computed in its original form (with TreeFARMS), and when RESPLIT is used as the Rashomon set generating algorithm. The reported Pearson correlation is computed between the top 2020 features. We see that it is very close to $1$, i.e. features that are important under RID will also remain important when RESPLIT is used. (bottom) The number of models across the bootstrapped Rashomon sets which split on a given feature. We note from the bar plots that RESPLIT is also able to generate a large number of trees - often times as many as TreeFARMS.
Parameters: $\lambda=0.02,\epsilon=0.01,\#$ bootstrapped datasets =10=10, depth budget =5=5, lookahead depth =3=3.
@@ A.6 Rashomon Importance Distribution Under RESPLIT: Quantile Binarization
In this section, we show similar results as in the previous section, but when datasets are binarized using feature quantiles. We chose $3$ quantiles per feature (corresponding to each 33rd of the distribution), resulting in datasets with $3\times$ the number of features. For most of these datasets, RID with TreeFARMS failed to run in reasonable time.
Figure 17: (top) Model Reliance under RESPLIT for the top 55 features. (bottom) The number of models across the bootstrapped Rashomon sets which split on a given feature. We note that the features which are important for RESPLIT under threshold guessing are also similarly important under quantile binarization, suggesting that our approach can generalize to different binarization schemes.
Parameters: $\lambda=0.02,\epsilon=0.01,\#$ bootstrapped datasets =10=10, depth budget =5=5, lookahead depth =3=3.
@@ A.7.1 Datasets
  • The Home Equity Line of Credit (HELOC) (heloc) dataset used for the Explainable ML Challenge. This dataset aims to predict the risk of loan default given the credit history of an individual. It consists of 23 features related to financial history, including FICO (credit) score, loan amount, number of delinquent accounts, credit inquiries, and other credit performance indicators. The dataset contains approximately 10,000 instances.
@@
  • Two recidivism datasets (COMPAS and Netherlands). COMPAS (LarsonMaKiAn16) aims to predict the likelihood of recidivism (reoffending) for individuals who have been arrested. The dataset consists of approximately $6000$ instances and includes $11$ features including demographic attributes, criminal history, risk of general recidivism, and chargesheet information. The Netherlands dataset (tollenaar2013method) is a similar recidivism dataset containing demographic and prior offense features for individuals, used to predict reoffending risk.
@@
  • The Covertype dataset (covertype), which aims to predict the forest cover (one of $7$ types) for areas of the Roosevelt National Forest in northern Colorado, based on cartographic data. It contains $54$ attributes derived from US Geological Survey data. These include continuous variables like elevation, aspect, slope, and others related to soil type and climate. The dataset has over 580,000 instances, each corresponding to a 3030m $\times$ 3030m patch of the forest.
@@
  • The Adult dataset (adult_2), which aims to predict whether an individual’s income exceeds $\mathdollar 50000$ per year based on demographic and occupational information. It contains around $50000$ train and test examples, with $14$ features.
@@
  • The Bike dataset (fanaee2013event), which contains a two-year historical log of bikeshare counts from 20112011-$2012$ in Washington D.C., USA. It contains features relating to the weather at every hour – with the aim being to predict the number of bike rentals in the city in that given time period.
@@
  • The Hypothyroid dataset (thyroid_disease_102), which contains medical records used to predict whether a patient has hypothyroidism based on thyroid function test results and other medical attributes. It includes categorical and continuous variables such as TSH (thyroid-stimulating hormone) levels, age, and presence of goiter, with thousands of instances.
@@
  • The Spambase dataset (spambase_94), which consists of email data used to classify messages as spam or not spam. The dataset contains 57 features extracted from email text, such as word frequencies, capital letter usage, and special character counts, with around 4600 instances.
@@
  • The Bank dataset (bank_marketing_222), which is used to predict whether a customer will subscribe to a bank term deposit based on features like age, job type, marital status, education level, and past marketing campaign success. It consists of approximately 4500 instances with 16 attributes.
@@
  • The HIV dataset (hiv) contains RNA samples from $2$ patients. The labels correspond to whether the observed HIV viral load is high or not.